# The Orbits of Isolated Dwarfs in the Local Group from New 3D Kinematics: Constraints on First Infall, Backsplash, and Quenching Mechanisms

### v1

@@ Abstract
software: Numpy (Harris et al., 2020), SciPy (Virtanen et al., 2020), Matplotlib (Hunter, 2007), IPython (Perez & Granger, 2007), Jupyter (Kluyver et al., 2016), Astropy (Astropy Collaboration et al., 2013; Price-Whelan et al., 2018; Astropy Collaboration et al., 2022), Gala (Price-Whelan, 2017)thanks: Patel and Bennet are co-first authors and have contributed
@@ 1 Introduction
The link between star formation and quenching is a key area of study in the dwarf galaxy regime, both in observations (Geha et al., 2012; Karunakaran et al., 2021, 2023; Font et al., 2022; Sales et al., 2022; Jones et al., 2023) and simulations (e.g. Fattahi et al., 2016; Wetzel et al., 2016; Simpson et al., 2018; Garrison-Kimmel et al., 2019; Libeskind et al., 2020; Samuel et al., 2022; Gutcke et al., 2022; Engler et al., 2023). In the Local Group (LG), most dwarf galaxies within the virial radius of the massive galaxies—Andromeda (M31) or the Milky Way (MW)—are quenched except for the most massive dwarfs, such as the Magellanic Clouds. On the other hand, isolated or field dwarfs—those outside the virial radius of a massive galaxy—are generally gas-rich and star forming (Spekkens et al., 2014; Putman et al., 2021), especially for dwarfs with stellar masses $M_{*}>10^{9}\,\rm M_{\odot}$ at more than 1.5 Mpc from a more massive galaxy (Geha et al., 2012). This trend is also seen in simulations of LG-like galaxy groups and isolated MW-like galaxies (e.g. Fattahi et al., 2016; Sawala et al., 2016; Wetzel et al., 2016; Simpson et al., 2018; Garrison-Kimmel et al., 2019; Libeskind et al., 2020; Engler et al., 2021, 2023; Font et al., 2021; Applebaum et al., 2021; Akins et al., 2021; Samuel et al., 2022).
@@
Isolated dwarfs are key to solving questions surrounding star formation and quenching, as it is difficult to disentangle the effects of internal (e.g., reionization and supernova feedback) and external (tidal and ram pressure stripping) processes on dwarf galaxies in the vicinity of massive galaxies (Gatto et al., 2013; Emerick et al., 2016; Tollerud & Peek, 2018; Putman et al., 2021). The influence of massive hosts can also extend to dwarfs far past their virial radius. Such galaxies are often referred to as backsplash galaxies. These are galaxies that are currently outside the virial radius of a massive host, but previously passed inside that radius and therefore may carry the imprints of any past interactions (e.g Gill et al., 2005; Sales et al., 2007; Fraternali et al., 2009; Santos-Santos et al., 2023). Such imprints include a lack of a significant HI gas reservoir, a lack of recent star formation, or morphological asymmetries.
@@
Distinguishing galaxies that are backsplash from those that are on first infall requires 3D position and velocity information (together referred to as 6D phase space information) to reconstruct plausible orbital histories for these objects. However, such 6D phase space information is challenging to obtain. Historically, almost all of our knowledge of LG galaxy orbits has come from line-of-sight (LOS) velocities ($V_{los}$). These motions are limited to motions towards or away from us, thus severely limiting characterization of wider dynamical properties and permitting a wide range of possible orbits. Proper motions (PMs), motions on the plane of the sky, provide the additional two components of motion needed for complete 6D characterization of a galaxy’s kinematics and position, and therefore more complex kinematic properties, e.g., orbits. With sufficiently accurate PMs, we can use these derived orbits to distinguish between first infall and backsplash galaxies (e.g., Bennet et al., 2024).
@@
In recent years PMs from Gaia have tightly constrained the 3D motions of M31 and M33 (van der Marel et al., 2019; Salomon et al., 2021a; Rusterucci et al., 2024), and have transformed our understanding of the MW satellites (Simon, 2018; Fritz et al., 2018a; Vasiliev, 2019; Patel et al., 2020; Battaglia et al., 2022a). However, motions of dwarfs outside of the outer MW halo are still mainly limited to one dimension. Gaia’s present precision (DR3) is not sufficient for measuring bulk PMs of low-mass systems beyond $\sim$400 kpc (Fritz et al., 2018a), and its magnitude limit at $G\sim 20.7$ mag means that some more distant objects will remain permanently out of reach, even with future data releases [^1] The current DR3 has a time baseline of 34 months, however even the final 10 year catalog (DR5) will not improve on this photometric limit..
@@
Efforts have been made to combine the Hubble Space Telescope (HST) and Gaia to get around some of these shortcomings (Massari et al., 2020; del Pino et al., 2022; Bennet et al., 2022; McKinnon et al., 2023; Warfield et al., 2023). However, these efforts are also limited to the areas where Gaia provides access to stellar populations that are bright enough and statistically significant, limiting it to nearby and star-forming objects, which often have upper main sequence member stars bright enough to appear in the Gaia catalog even beyond the MW halo. Efforts are also underway to combine Gaia with Euclid (Libralato et al., 2024), though this has the same issues as combining HST and Gaia.
@@
Meanwhile, HST can measure the PMs of even the most distant of LG galaxies, or even potentially beyond the LG, when two or more epochs of observations are available with sufficient temporal separation (time baseline) between them (e.g. Sohn et al., 2013, 2017, 2020). Here we report the PMs of six distant isolated LG dwarf galaxies using HST. Our sample members were chosen due to their isolated status and lack of constraining PMs in the literature, with only two of the six having any measurements, albeit with highly-uncertain tangential velocities (McConnachie & Venn, 2020a, b; Battaglia et al., 2022b). These two galaxies have also been examined through a combination of Gaia and HST (Bennet et al., 2024, hereafter B24), and while those uncertainties improved on the Gaia-only results, they are still not constrained sufficiently to provide physically-meaningful orbits.
@@ 2 Data
This work uses data from the Hubble Space Telescope (HST) to derive PMs for the six dwarfs in our sample, which are discussed in Section 2.3 and listed in Table 2. Additionally, we have obtained Green Bank Telescope (GBT) observations of the HI distribution in Cetus. The latter improves upon previous observations from Putman et al. (2021), and will be further discussed in Section 4.2 when we relate our orbital results to Cetus’s SFH and gas content.
@@ 2.2 Green Bank Telescopes observations
Of our dwarf sample, only Cetus and Tucana do not have well-established literature HI masses (Putman et al., 2021). Measurements of HI masses will be discussed in Section 4, as they provide insight into the environment these dwarfs have survived in. To potentially detect HI in these dwarfs, we observed them with the GBT concurrently with this analysis. Tucana is a southern-sky object and not visible from the GBT, so only Cetus was observed.
@@ 2.3 Proper Motion Analysis
Note. — $\mu^{*}_{\alpha}$ = $\mu_{\alpha}\times cos(\delta),$ where $\mu_{\alpha}$ is the time derivative of the RA in units of angle (not hour, minute, second units). These are on-sky values and are not corrected for solar reflex motion. $V_{\rm trans,\alpha}$ and $V_{\rm trans,\delta}$ refer to the transverse velocities computed using the proper motions and distances provided. Uncertainties on $V_{\rm trans,\alpha}$ and $V_{\rm trans,\delta}$ do not include the propagation of distance uncertainties. References for the literature heliocentric distances, column 2: Aquarius (Ordoñez & Sarajedini, 2016), Cetus & Tucana (Dambis et al., 2013), Leo T (Clementini et al., 2012), Pegasus & Pisces (Savino et al., 2025). References for the literature LOS velocities, column 4: Aquarius (Springob et al., 2005), Cetus (Kashibadze & Karachentsev, 2018), Leo T (Haynes et al., 2018), Pegasus & Pisces (Huchtmeier et al., 2003), Tucana (Fraternali et al., 2009).
To measure the PMs of our dwarf sample, we closely follow the methods used extensively in our previous work (Sohn et al., 2012, 2013, 2017, 2020; Bennet et al., 2024). We will summarize our methodology here, but invite interested readers to read these other works for more details.
We started by downloading the fully pipeline-processed image for each exposure. Specifically, we downloaded the _flc.fits images, which not only include all standard CCD reduction steps (bias and dark subtraction, flat-fielding, etc.), but also have been processed for the imperfect charge transfer efficiency (CTE) using the correction algorithms of Anderson & Bedin (2010). We determine a position and flux for each point source from the _flc.fits images using the FORTRAN code hst1pass (Anderson, 2022). We applied corrections to the positions using the methodology from (Anderson & King, 2006), with updated ACS/WFC geometric distortions based on Kozhurina-Platais et al. (2015); these were further extended to include time-dependent distortion variations beyond 2020 (V. Kozhurina-Platais, private communication). For each dwarf galaxy, we create high-resolution stacked images with pixel scales of $0.025$ mas using images from the reference epoch. We use the epoch with the most exposures as the reference epoch across our sample (see Table 1 for exposure counts for each epoch).
@@
We then construct an empirical “template” for each point source and background galaxy by supersampling the objects extracted from the high-resolution stack of images. Supersampling uses the fact that the dithers between HST exposures are a non-integer number of pixels; thus, stacked images are super-sampled by a factor of two relative to the native ACS/ WFC pixel scale for better spatial resolution. Each template takes into account the point-spread function (PSF), the morphology of background galaxies, and the pixel binning (Sohn et al., 2013). These templates are fitted on each individual _flc.fits image for measuring the position of each point source or galaxy. We fit the templates directly onto the individual images for the reference epoch since the templates were constructed from that epoch. For the other non-reference epoch(s), we employ $7\times 7$ pixel convolution kernels when fitting templates to allow for differences in PSF between the different epochs. These convolution kernels were derived by comparing PSFs of numerous bright and isolated point sources associated with the target dwarf galaxies between the reference and other epochs (Sohn et al., 2013). At this stage, we are left with template-fitted positions of all point sources and background galaxies in all individual images in each epoch.
@@
Once the reflex displacements of background galaxies are measured for each non-reference epoch, we calculate the final PMs. For those dwarfs with only two epochs of observations, i.e., Aquarius and Pegasus, the PM calculations were carried out straightforwardly following those in Sohn et al. (2020). In summary, we take the error-weighted average of all displacements of background galaxies with respect to dwarf member stars for each individual first-epoch exposure to obtain an independent PM estimate for each exposure. The PM was obtained from the error-weighted mean weighted by the inverse variance of individual PM exposure estimates. Similarly, associated uncertainties in the mean were obtained by summing the inverse variances to get the inverse variance of the mean, also from the individual exposure PM estimates.
@@
Note. — Three-dimensional position and velocity vectors derived, in Galactocentric coordinates, from the galaxy positions, distances, LOS velocities, and PM measurements as reported in Table 2. The Galactocentric distance of the Sun and the circular velocity of the local standard of rest (LSR) are adopted from McMillan (2011). Solar peculiar velocities with respect to the LSR are taken from Schönrich et al. (2010).
@@ 3.1 Galactocentric Coordinates
Figure 1 shows the magnitude of the 3D velocity for each dwarf relative to the distance from the MW (middle panel) and M31 (right panel) compared to escape velocity curves for host galaxy halos of masses between $1-4\times 10^{12}\,\rm M_{\odot}$. This mass range captures the literature mass estimates for both the MW and M31 (see reviews for MW and M31, respectively; Wang et al., 2020; Bhattacharya, 2023). The relative distances and velocities are computed assuming the distance and PM values for M31 listed in Section 3.2.
Escape velocity curves were computed assuming NFW halo profiles. This work uses a virial mass of $1\times 10^{12}\,\rm M_{\odot}$ for the MW and $2\times 10^{12}\,\rm M_{\odot}$ for M31. However, only four of the six dwarfs are bound to the MW assuming $\rm M_{vir}=1\times 10^{12}\,\rm M_{\odot}$ (Patel et al., 2018). For M31 ($\rm M_{vir}=2\times 10^{12}\,\rm M_{\odot}$; Fardal et al., 2013; Villanueva-Domingo et al., 2022, e.g.,), only two of six dwarfs are bound to M31. While not all dwarfs are bound to either the MW or M31, they are bound to the joint potential of the MW+M31.
@@ 3.2 Orbit Methodology
To reconstruct orbital histories for the six LG dwarfs in our sample, we follow the methodology of B24. In brief, the 6D phase space coordinates reported in Table 3 are used to initialize backward numerical orbit integration for 6 Gyr. We also compute orbits 6 Gyr into the future. Our chosen integration period of 6 Gyr corresponds to the time by which the MW and M31 had assembled about 80% of their mass (Santistevan et al., 2020). For each galaxy, we compute a 5-body orbit accounting for the gravitational potential of the MW, M31, LMC [^3] Note that we do not correct for the perturbation introduced by the LMC, which offsets the outer MW halo from the inner MW halo (Garavito-Camargo et al., 2019). As discussed in 3, none of our six dwarfs closely approach the MW or the LMC, so this effect is likely negligible for the purposes of this work., M33, and each dwarf galaxy of interest. For the four massive LG galaxies, we use identical masses and parameters as those in B24, which are originally adopted from Patel et al. (2017a). Each galaxy is represented as an extended mass distribution. The MW and M31 are represented as three-component potentials with an NFW halo, Miyamoto-Nagai disk, and a Hernquist bulge. Dark matter halos are adiabatically contracted using the CONTRA code (Gnedin et al., 2004). Both the LMC and M33 are modeled as single-component Plummer spheres. Assumed total masses for each galaxy are as follows: $(2,1,0.25,0.18)\times 10^{12}\,\rm M_{\odot}$ for M31, the MW, M33, and the LMC, respectively.
@@
We use M31 and M33 distances from Savino et al. (776.2 kpc and 859 kpc, respectively; 2022), M31’s PM from Salomon et al. (2021b), and M33’s PM from Brunthaler et al. (2005). 6D phase space properties for the LMC are identical to those in Patel et al. (2017a). We encourage interested readers to see B24 and Patel et al. (2017a) for model parameters and details relevant to the orbital integration scheme.
Orbits initialized with the mean 6D phase space coordinates listed in Table 3 will be referred to as direct orbital histories, as in past work (Sohn et al., 2020; Patel et al., 2020; Bennet et al., 2024). Uncertainties on orbital histories will be discussed in Section 3.4.

### v2

@@ Abstract
software: Numpy (C. R. Harris et al., 2020), SciPy (P. Virtanen et al., 2020), Matplotlib (J. D. Hunter, 2007), IPython (F. Perez & B. E. Granger, 2007), Jupyter (T. Kluyver et al., 2016), Astropy (Astropy Collaboration et al., 2013; A. M. Price-Whelan et al., 2018; Astropy Collaboration et al., 2022), Gala (A. M. Price-Whelan, 2017)thanks: Patel and Bennet are co-first authors and have contributed
@@ 1 Introduction
The link between star formation and quenching is a key area of study in the dwarf galaxy regime, both in observations (M. Geha et al., 2012; A. Karunakaran et al., 2021, 2023; A. S. Font et al., 2022; L. V. Sales et al., 2022; M. G. Jones et al., 2023) and simulations (e.g. A. Fattahi et al., 2016; A. R. Wetzel et al., 2016; C. M. Simpson et al., 2018; S. Garrison-Kimmel et al., 2019; N. I. Libeskind et al., 2020; J. Samuel et al., 2022; T. A. Gutcke et al., 2022; C. Engler et al., 2023). In the Local Group (LG), most dwarf galaxies within the virial radius of the massive galaxies—Andromeda (M31) or the Milky Way (MW)—are quenched except for the most massive dwarfs, such as the Magellanic Clouds. On the other hand, isolated or field dwarfs—those outside the virial radius of a massive galaxy—are generally gas-rich and star forming (K. Spekkens et al., 2014; M. E. Putman et al., 2021), especially for dwarfs with stellar masses $M_{*}>10^{9}\,\rm M_{\odot}$ at more than 1.5 Mpc from a more massive galaxy (M. Geha et al., 2012). This trend is also seen in simulations of LG-like galaxy groups and isolated MW-like galaxies (e.g. A. Fattahi et al., 2016; T. Sawala et al., 2016; A. R. Wetzel et al., 2016; C. M. Simpson et al., 2018; S. Garrison-Kimmel et al., 2019; N. I. Libeskind et al., 2020; C. Engler et al., 2021, 2023; A. S. Font et al., 2021; E. Applebaum et al., 2021; H. B. Akins et al., 2021; J. Samuel et al., 2022).
@@
Isolated dwarfs are key to solving questions surrounding star formation and quenching, as it is difficult to disentangle the effects of internal (e.g., reionization and supernova feedback) and external (tidal and ram pressure stripping) processes on dwarf galaxies in the vicinity of massive galaxies (A. Gatto et al., 2013; A. Emerick et al., 2016; E. J. Tollerud & J. E. G. Peek, 2018; M. E. Putman et al., 2021). The influence of massive hosts can also extend to dwarfs far past their virial radius. Such galaxies are often referred to as backsplash galaxies. These are galaxies that are currently outside the virial radius of a massive host, but previously passed inside that radius and therefore may carry the imprints of any past interactions (e.g S. P. D. Gill et al., 2005; L. V. Sales et al., 2007; F. Fraternali et al., 2009; I. M. E. Santos-Santos et al., 2023). Such imprints include a lack of a significant HI gas reservoir, a lack of recent star formation, or morphological asymmetries.
@@
Distinguishing galaxies that are backsplash from those that are on first infall requires 3D position and velocity information (together referred to as 6D phase space information) to reconstruct plausible orbital histories for these objects. However, such 6D phase space information is challenging to obtain. Historically, almost all of our knowledge of LG galaxy orbits has come from line-of-sight (LOS) velocities ($V_{los}$). These motions are limited to motions towards or away from us, thus severely limiting characterization of wider dynamical properties and permitting a wide range of possible orbits. Proper motions (PMs), motions on the plane of the sky, provide the additional two components of motion needed for complete 6D characterization of a galaxy’s kinematics and position, and therefore more complex kinematic properties, e.g., orbits. With sufficiently accurate PMs, we can use these derived orbits to distinguish between first infall and backsplash galaxies (e.g., P. Bennet et al., 2024).
@@
In recent years PMs from Gaia have tightly constrained the 3D motions of M31 and M33 (R. P. van der Marel et al., 2019; J. B. Salomon et al., 2021a; S. Rusterucci et al., 2024), and have transformed our understanding of the MW satellites (J. D. Simon, 2018; T. K. Fritz et al., 2018a; E. Vasiliev, 2019; E. Patel et al., 2020; G. Battaglia et al., 2022a). However, motions of dwarfs outside of the outer MW halo are still mainly limited to one dimension. Gaia’s present precision (DR3) is not sufficient for measuring bulk PMs of low-mass systems beyond $\sim$400 kpc (T. K. Fritz et al., 2018a), and its magnitude limit at $G\sim 20.7$ mag means that some more distant objects will remain permanently out of reach, even with future data releases [^1] The current DR3 has a time baseline of 34 months, however even the final 10 year catalog (DR5) will not improve on this photometric limit..
@@
Efforts have been made to combine the Hubble Space Telescope (HST) and Gaia to get around some of these shortcomings (D. Massari et al., 2020; A. del Pino et al., 2022; P. Bennet et al., 2022; K. A. McKinnon et al., 2023; J. T. Warfield et al., 2023). However, these efforts are also limited to the areas where Gaia provides access to stellar populations that are bright enough and statistically significant, limiting it to nearby and star-forming objects, which often have upper main sequence member stars bright enough to appear in the Gaia catalog even beyond the MW halo. Efforts are also underway to combine Gaia with Euclid (M. Libralato et al., 2024), though this has the same issues as combining HST and Gaia.
@@
Meanwhile, HST can measure the PMs of even the most distant of LG galaxies, or even potentially beyond the LG, when two or more epochs of observations are available with sufficient temporal separation (time baseline) between them (e.g. S. T. Sohn et al., 2013, 2017, 2020). Here we report the PMs of six distant isolated LG dwarf galaxies using HST. Our sample members were chosen due to their isolated status and lack of constraining PMs in the literature, with only two of the six having any measurements, albeit with highly-uncertain tangential velocities (A. W. McConnachie & K. A. Venn, 2020a, b; G. Battaglia et al., 2022b). These two galaxies have also been examined through a combination of Gaia and HST (P. Bennet et al., 2024, hereafter B24), and while those uncertainties improved on the Gaia-only results, they are still not constrained sufficiently to provide physically-meaningful orbits.
@@ 2 Data
This work uses data from the Hubble Space Telescope (HST) to derive PMs for the six dwarfs in our sample, which are discussed in Section 2.3 and listed in Table 2. Additionally, we have obtained Green Bank Telescope (GBT) observations of the HI distribution in Cetus. The latter improves upon previous observations from M. E. Putman et al. (2021), and will be further discussed in Section 4.2 when we relate our orbital results to Cetus’s SFH and gas content.
@@ 2.2 Green Bank Telescopes observations
Of our dwarf sample, only Cetus and Tucana do not have well-established literature HI masses (M. E. Putman et al., 2021). Measurements of HI masses will be discussed in Section 4, as they provide insight into the environment these dwarfs have survived in. To potentially detect HI in these dwarfs, we observed them with the GBT concurrently with this analysis. Tucana is a southern-sky object and not visible from the GBT, so only Cetus was observed.
@@
Note. — $\mu^{*}_{\alpha}$ = $\mu_{\alpha}\times cos(\delta),$ where $\mu_{\alpha}$ is the time derivative of the RA in units of angle (not hour, minute, second units). These are on-sky values and are not corrected for solar reflex motion. $V_{\rm trans,\alpha}$ and $V_{\rm trans,\delta}$ refer to the transverse velocities computed using the proper motions and distances provided. Uncertainties on $V_{\rm trans,\alpha}$ and $V_{\rm trans,\delta}$ do not include the propagation of distance uncertainties. References for the literature heliocentric distances, column 2: Aquarius (A. J. Ordoñez & A. Sarajedini, 2016), Cetus & Tucana (A. K. Dambis et al., 2013), Leo T (G. Clementini et al., 2012), Pegasus & Pisces (A. Savino et al., 2025). References for the literature LOS velocities, column 4: Aquarius (C. M. Springob et al., 2005), Cetus (O. G. Kashibadze & I. D. Karachentsev, 2018), Leo T (M. P. Haynes et al., 2018), Pegasus & Pisces (W. K. Huchtmeier et al., 2003), Tucana (F. Fraternali et al., 2009).
2.3 Proper Motion Analysis
To measure the PMs of our dwarf sample, we closely follow the methods used extensively in our previous work (S. T. Sohn et al., 2012, 2013, 2017, 2020; P. Bennet et al., 2024). We will summarize our methodology here, but invite interested readers to read these other works for more details.
We started by downloading the fully pipeline-processed image for each exposure. Specifically, we downloaded the _flc.fits images, which not only include all standard CCD reduction steps (bias and dark subtraction, flat-fielding, etc.), but also have been processed for the imperfect charge transfer efficiency (CTE) using the correction algorithms of J. Anderson & L. R. Bedin (2010). We determine a position and flux for each point source from the _flc.fits images using the FORTRAN code hst1pass (J. Anderson, 2022). We applied corrections to the positions using the methodology from (J. Anderson & I. R. King, 2006), with updated ACS/WFC geometric distortions based on V. Kozhurina-Platais et al. (2015); these were further extended to include time-dependent distortion variations beyond 2020 (V. Kozhurina-Platais, private communication). For each dwarf galaxy, we create high-resolution stacked images with pixel scales of $0.025$ mas using images from the reference epoch. We use the epoch with the most exposures as the reference epoch across our sample (see Table 1 for exposure counts for each epoch).
@@ 2.3 Proper Motion Analysis
We then construct an empirical “template” for each point source and background galaxy by supersampling the objects extracted from the high-resolution stack of images. Supersampling uses the fact that the dithers between HST exposures are a non-integer number of pixels; thus, stacked images are super-sampled by a factor of two relative to the native ACS/ WFC pixel scale for better spatial resolution. Each template takes into account the point-spread function (PSF), the morphology of background galaxies, and the pixel binning (S. T. Sohn et al., 2013). These templates are fitted on each individual _flc.fits image for measuring the position of each point source or galaxy. We fit the templates directly onto the individual images for the reference epoch since the templates were constructed from that epoch. For the other non-reference epoch(s), we employ $7\times 7$ pixel convolution kernels when fitting templates to allow for differences in PSF between the different epochs. These convolution kernels were derived by comparing PSFs of numerous bright and isolated point sources associated with the target dwarf galaxies between the reference and other epochs (S. T. Sohn et al., 2013). At this stage, we are left with template-fitted positions of all point sources and background galaxies in all individual images in each epoch.
@@
Once the reflex displacements of background galaxies are measured for each non-reference epoch, we calculate the final PMs. For those dwarfs with only two epochs of observations, i.e., Aquarius and Pegasus, the PM calculations were carried out straightforwardly following those in S. T. Sohn et al. (2020). In summary, we take the error-weighted average of all displacements of background galaxies with respect to dwarf member stars for each individual first-epoch exposure to obtain an independent PM estimate for each exposure. The PM was obtained from the error-weighted mean weighted by the inverse variance of individual PM exposure estimates. Similarly, associated uncertainties in the mean were obtained by summing the inverse variances to get the inverse variance of the mean, also from the individual exposure PM estimates.
@@
Note. — Three-dimensional position and velocity vectors derived, in Galactocentric coordinates, from the galaxy positions, distances, LOS velocities, and PM measurements as reported in Table 2. The Galactocentric distance of the Sun and the circular velocity of the local standard of rest (LSR) are adopted from P. J. McMillan (2011). Solar peculiar velocities with respect to the LSR are taken from R. Schönrich et al. (2010).
@@ 3.1 Galactocentric Coordinates
Figure 1 shows the magnitude of the 3D velocity for each dwarf relative to the distance from the MW (middle panel) and M31 (right panel) compared to escape velocity curves for host galaxy halos of masses between $1-4\times 10^{12}\,\rm M_{\odot}$. This mass range captures the literature mass estimates for both the MW and M31 (see reviews for MW and M31, respectively; W. Wang et al., 2020; S. Bhattacharya, 2023). The relative distances and velocities are computed assuming the distance and PM values for M31 listed in Section 3.2.
Escape velocity curves were computed assuming NFW halo profiles. This work uses a virial mass of $1\times 10^{12}\,\rm M_{\odot}$ for the MW and $2\times 10^{12}\,\rm M_{\odot}$ for M31. However, only four of the six dwarfs are bound to the MW assuming $\rm M_{vir}=1\times 10^{12}\,\rm M_{\odot}$ (E. Patel et al., 2018). For M31 ($\rm M_{vir}=2\times 10^{12}\,\rm M_{\odot}$; M. A. Fardal et al., 2013; P. Villanueva-Domingo et al., 2022, e.g.,), only two of six dwarfs are bound to M31. While not all dwarfs are bound to either the MW or M31, they are bound to the joint potential of the MW+M31.
@@ 3.2 Orbit Methodology
To reconstruct orbital histories for the six LG dwarfs in our sample, we follow the methodology of B24. In brief, the 6D phase space coordinates reported in Table 3 are used to initialize backward numerical orbit integration for 6 Gyr. We also compute orbits 6 Gyr into the future. Our chosen integration period of 6 Gyr corresponds to the time by which the MW and M31 had assembled about 80% of their mass (I. B. Santistevan et al., 2020). For each galaxy, we compute a 5-body orbit accounting for the gravitational potential of the MW, M31, LMC [^3] Note that we do not correct for the perturbation introduced by the LMC, which offsets the outer MW halo from the inner MW halo (N. Garavito-Camargo et al., 2019). As discussed in 3, none of our six dwarfs closely approach the MW or the LMC, so this effect is likely negligible for the purposes of this work., M33, and each dwarf galaxy of interest. For the four massive LG galaxies, we use identical masses and parameters as those in B24, which are originally adopted from E. Patel et al. (2017a). Each galaxy is represented as an extended mass distribution. The MW and M31 are represented as three-component potentials with an NFW halo, Miyamoto-Nagai disk, and a Hernquist bulge. Dark matter halos are adiabatically contracted using the CONTRA code (O. Y. Gnedin et al., 2004). Both the LMC and M33 are modeled as single-component Plummer spheres. Assumed total masses for each galaxy are as follows: $(2,1,0.25,0.18)\times 10^{12}\,\rm M_{\odot}$ for M31, the MW, M33, and the LMC, respectively.
@@
We use M31 and M33 distances from A. Savino et al. (776.2 kpc and 859 kpc, respectively; 2022), M31’s PM from J. B. Salomon et al. (2021b), and M33’s PM from A. Brunthaler et al. (2005). 6D phase space properties for the LMC are identical to those in E. Patel et al. (2017a). We encourage interested readers to see B24 and E. Patel et al. (2017a) for model parameters and details relevant to the orbital integration scheme.
Orbits initialized with the mean 6D phase space coordinates listed in Table 3 will be referred to as direct orbital histories, as in past work (S. T. Sohn et al., 2020; E. Patel et al., 2020; P. Bennet et al., 2024). Uncertainties on orbital histories will be discussed in Section 3.4.